# The Artificial Scientist: in-transit Machine Learning of Plasma Simulations

### v2

@@ VI Conclusion
By training a machine learning model on-the-fly on a PIConGPU simulation of the Kelvin-Helmholtz instability, where data is streamed via openPMD with ADIOS2, we demonstrate a realization of the Artificial Scientist solving an ill-posed inverse problem from plasma physics. This allows us to circumvent the lack of adequate disk capacity and bandwidth, often a bottleneck in large-scale systems, by enabling data to reside in-memory and be distributed via network between the nodes. Scientifically, this work also demonstrates as a proof of principle, that it is possible to learn correlations from a physics simulation on-the-fly. On full Frontier, we demonstrate streaming at scale using openPMD as well as performance studies at scale for general plasma physics.
@@ Acknowledgments
This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
The research was supported by Laserlab-Europe EU-H2020 GA no. 871124.
The support work from the ADIOS team was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program, under the “RAPIDS Institute”.

### v3

@@ VI Conclusion
By training a machine learning model on-the-fly on a PIConGPU simulation of the Kelvin-Helmholtz instability, where data is streamed via openPMD with ADIOS2, we demonstrate a realization of the Artificial Scientist solving an ill-posed inverse problem from plasma physics. This allows us to circumvent the lack of adequate disk capacity and bandwidth, often a bottleneck in large-scale systems, by enabling data to reside in-memory and be distributed via network between the nodes. Scientifically, as a proof of principle, this work also demonstrates that it is possible to learn correlations from a physics simulation on-the-fly. The presented workflow is scalable from local clusters to Top-1 supercomputers and is tested across GPU vendors. This is possible due to PIConGPU’s foundation on the abstraction layer alpaka and due to the availability of various ports of PyTorch. On full Frontier, we demonstrate streaming at scale using openPMD as well as performance studies at scale for general plasma physics.
@@ Acknowledgments
The authors thank Patrick Stiller and Nico Hoffmann for fruitful discussions. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. This research was supported by the Helmholtz Association Initiative and Networking Fund in the frame of Helmholtz AI, by Laserlab-Europe EU-H2020 GA no. 871124 and the European Union through Grant Agreement No. 101093261 (Plasma-PEPSC). The support work from the ADIOS team was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program, under the “RAPIDS Institute”.